# Grounding-IQA: Multimodal Language Grounding Model for Image Quality Assessment

### v2

@@ Abstract
The development of multimodal large language models (MLLMs) enables the evaluation of image quality through natural language descriptions. This advancement allows for more detailed assessments. However, these MLLM-based IQA methods primarily rely on general contextual descriptions, sometimes limiting fine-grained quality assessment. To address this limitation, we introduce a new image quality assessment (IQA) task paradigm, grounding-IQA. This paradigm integrates multimodal referring and grounding with IQA to realize more fine-grained quality perception. Specifically, grounding-IQA comprises two subtasks: grounding-IQA-description (GIQA-DES) and visual question answering (GIQA-VQA). GIQA-DES involves detailed descriptions with precise locations (e.g., bounding boxes), while GIQA-VQA focuses on quality QA for local regions. To realize grounding-IQA, we construct a corresponding dataset, GIQA-160K, through our proposed automated annotation pipeline. Furthermore, we develop a well-designed benchmark, GIQA-Bench. The benchmark comprehensively evaluates the model grounding-IQA performance from three perspectives: description quality, VQA accuracy, and grounding precision. Experiments demonstrate that our proposed task paradigm, dataset, and benchmark facilitate the more fine-grained IQA application. Code: https://github.com/zhengchen1999/Grounding-IQA.
@@
1 Introduction
Image quality assessment (IQA) seeks to evaluate image quality in alignment with human perception. As a fundamental task in low-level vision, IQA is critical across multiple fields, e.g., image processing [57, 19], media transmission [50], and generative artificial intelligence [16]. However, this task is challenging since the human visual system is inherently subjective and complex to model [41]. To enhance evaluation precision, substantial research efforts continue to be dedicated to this area [29, 5, 39, 43].
Traditional IQA methods employ handcrafted metrics to estimate quality scores [41, 30]. With advancements in deep neural networks, learning specific priors from large datasets enables more accurate score predictions [11, 1, 57, 10, 13]. Nevertheless, score-based IQA methods face challenges in complex scenarios. In such cases, image quality is influenced by multiple factors that a single score cannot effectively express [53]. Recently, the emergence of multimodal large language models (MLLMs) [21, 33, 49] with strong visual and linguistic perception capabilities provides an alternative to score-based IQA [42, 46]. These models achieve more detailed and accurate image assessments through description and reasoning. However, current MLLM-based IQA methods [43, 52] primarily rely on general contextual descriptions, which sometimes limits fine-grained quality assessments. For instance, in Fig. 2a, the existing method (i.e., Q-Instruct [43]) describes the objects/areas affecting image quality through language, but cannot provide precise location information. Moreover, in Fig. 2b, for local perception, the language referring may not accurately pinpoint the target. These limitations restrict the application of MLLMs in comprehensive low-level perception and understanding, especially for fine-grained cases.
@@ 1 Introduction
To address these challenges and unleash the potential of MLLMs in fine-grained image quality understanding, we introduce grounding-IQA. This is a novel IQA task paradigm that integrates multimodal referring (position in) and grounding (position out[28, 4, 33] with image quality assessment. Specifically, we categorize grounding-IQA into two sub-tasks: (1) Grounding-IQA-Description (GIQA-DES). As illustrated in Fig. 2a, this task requires generating descriptive assessments of image quality while providing precise locations (i.e., bounding boxes) for important objects/regions impacting quality. (2) Grounding-IQA-Visual Question Answering (GIQA-VQA). As shown in Fig. 2b, this task involves QA about low-level attributes of images, especially regarding local objects. It includes addressing questions with specific coordinates (referring) or providing answers with precise positions (grounding).
@@
Since existing datasets can not realize grounding-IQA well [21, 51, 43], we construct a new dataset, GIQA-160K, based on the proposed paradigm. This dataset can enhance the grounding-IQA capabilities of current MLLMs. The dataset comprises 160K instruction-tuning data with 40K images from diverse domains. Specifically, the dataset corresponds to two sub-tasks: GIQA-DES includes 60K corresponding data, and GIQA-VQA contains 100K related data. To construct the corresponding dataset, we design an automated annotation pipeline. The automated pipeline generates the GIQA-160K through the public IQA dataset [43, 52] (with the human-annotated description). (1) For GIQA-DES. The task includes detailed descriptions with coordinates. We generate the data through advanced vision [24] and language [6] models. Through these models, we extract and filter objects and corresponding coordinates from existing descriptions and images. Meanwhile, coordinates are expressed in natural language and attached to text. This avoids extra specialized tokens and ensures data compatibility. (2) For GIQA-VQA. Inspired by previous work [43, 51, 18], we construct the required data from the detailed descriptions in GIQA-DES via the LLM. We use specific QA templates (i.e., “Yes/No”, abbreviated as Y; “What/How/Why”, abbreviated as W) and emphasize location-specific objects to generate appropriate data. The coordinates are also combined with the generated QA.
@@
Fine-tuning on the GIQA-160K dataset enables existing pre-trained MLLMs to achieve impressive grounding-IQA capabilities. As shown in Fig. 2, the fine-tuned model can ground key objects affecting image quality, and perform more fine-grained assessments based on reference coordinates. Moreover, to comprehensively evaluate the model performance on the grounding-IQA task, we propose a well-designed benchmark, GIQA-Bench. This benchmark includes 100 varying types and quality images, corresponding to 100 GIQA-Des and 150 GIQA-VQA test samples. Each sample is annotated over multiple rounds by at least three experts. We quantitatively assess grounding-IQA performance in three aspects: (1) assessment description quality (i.e., BLEU@4, LLM-Score); (2) VQA accuracy (i.e., Accuracy); and (3) grounding precision (i.e., mIoU, Tag-Recall). We test recent MLLMs, with results shown in Fig. 1. Observations indicate significant improvement in grounding-IQA after fine-tuning with GIQA-160K.
Overall, our contributions are threefold:
@@ 2.1 Image Quality Assessment
Score-based Methods. Most current IQA methods are score-based. Early IQA approaches compute scores through handcrafted image data metrics [41, 55, 31, 29, 36]. For instance, PSNR calculates the ratio of signal to noise. NIQE [30] relies on the statistical characteristics of natural images. However, these methods show a gap in quality perception compared to human judgment and are unsuitable for complex scenarios. With the development of the deep neural network, learning-based IQA methods have gradually become mainstream [11, 25, 48, 2, 37, 40]. These methods leverage data-driven training to achieve more accurate quality assessments. For example, LPIPS [57] applies the convolutional neural network to compute scores. MUSIQ [13] employs the Transformer to extract multi-scale features for score prediction. Moreover, meta-learning [62], multimodal models [39, 59], and graph neural networks [38] have been adopted to further improve IQA performance. However, score-based IQA methods face limitations in complex scenarios. The simple score cannot effectively represent the multiple aspects affecting image quality.
MLLM-based Methods. Multimodal large language models (MLLMs) exhibit remarkable multimodal (language/vision) understanding by integrating visual modules into LLMs [21, 61, 56, 17, 49]. MLLMs achieve outstanding performance in various multimodal tasks, including visual question answering and image captioning. Recently, several studies have also demonstrated the potential of MLLMs in low-level visual perception and assessment [43, 52, 52, 45, 3]. For instance, Q-Instruct [43] constructs a multimodal dataset to enhance. Q-Align [44] guides MLLMs in scoring by defining discrete text-based levels. DepictQA [53] enables quality comparison and reasoning based on reference images. These approaches advance the application of MLLMs in IQA, achieving more accurate assessments. Nevertheless, these models primarily rely on contextual descriptions, and face limitations in fine-grained applications, e.g., local perception.
@@ 2.2 Multimodal Referring and Grounding
Multimodal spatial perception involves referring and grounding. Referring requires the model to understand the specific region based on position input, e.g., region-level captioning [14, 54]. Grounding, on the other hand, involves the model describing the region by outputting position, e.g., referring expression comprehension [12, 27]. Currently, MLLMs perform impressively in spatial perception, further advancing these tasks. Some methods focus on grounding, achieving complex reasoning [15] or multi-object [35] segmentation. Meanwhile, other approaches, e.g., GPT4RoI [58], emphasize understanding specific regions (referring). Furthermore, some works unify referring and grounding [4, 18, 34]. Kosmos-2 [33] utilizes bounding box coordinates to integrate both aspects. Ferret [51] extends to referring to arbitrary shapes. Additionally, in IQA, Q-Ground [3] achieves degradation region grounding but lacks referring capabilities. Overall, compared to previous work, our Grounding-IQA integrates multimodal referring and grounding with IQA. This new IQA paradigm enables more fine-grained and flexible quality perception.
Figure 3: The illustration of the automated annotation pipeline. (a) GIQA-DES Pipeline: Constructs the $\bm{answer}$ from the given image and corresponding description through a four-stage process, while the $\bm{question}$ comes from a predefined question pool. (b) GIQA-VQA Pipeline: Generates the corresponding QA data utilizing descriptions from GIQA-DES and the advanced LLM, Llama3 [6].
@@ 3.1 Grounding-IQA
GIQA-DES. The task requires the model to provide a detailed description of image quality. Additionally, it needs accurate location information (e.g., bounding box) for key objects/regions that impact image quality, as shown in Fig. 5a. This corresponds to the fact that humans consider not only the overall quality (e.g., image clarity) but also the quality of specific objects or locations when assessing image quality. Meanwhile, accurate location information also enables targeted information for downstream tasks (e.g., image editing). This task is similar to grounded image captioning [60], but places greater emphasis on low-level attributes. While some MLLMs [4, 33, 18] perform well in grounded image captioning, they still struggle with quality perception. We demonstrated it in Sec. 4.3.
GIQA-VQA. The second task focuses on the question-answering ability in low-level perception, particularly for local objects. Corresponding to multimodal referring and grounding, this task can be divided into two scenarios. Referring: querying low-level attributes in the specified region (input position), as shown in Fig. 5b. Grounding: providing answers that include specific locations (output position) based on the question, as depicted in Fig. 5b. These two scenarios are related to region captioning [60] and phrase grounding [60], respectively. However, like GIQA-DES, GIQA-VQA involves quality perception, which is challenging for current MLLMs. We evaluate in Sec. 4.3.
@@ 3.2 Automated Annotation Pipeline
Data is essential for achieving Grounding-IQA. Therefore, we construct an automated annotation pipeline to generate data (i.e., GIQA-160K). This pipeline leverages public IQA datasets [43, 52] that contain human-annotated descriptions. Following previous schemes [21, 49], the data format is $\bm{image}$, $\bm{question}$, $\bm{answer}$}. The $\bm{image}$ is the evaluation target. Depending on the sub-task, the $\bm{question}$ and $\bm{answer}$ fields may include precise coordinates (i.e., bounding box), in addition to text. We introduce the pipeline below; refer to the supplement for more details.
For GIQA-DES. In this task, the $\bm{question}$ is relatively fixed, as in Fig. 5a. To enhance data diversity, we construct a pool of 15 similar questions. For each data, the $\bm{question}$ is randomly selected from the pool. For the $\bm{answer}$, it is a detailed description with coordinates. We construct it via a four-stage process from existing images and associated description, as illustrated in Fig. 3: (1) Stage-1: object tag extraction; (2) Stage-2: bounding box detection; (3) Stage-3: box refinement (filter and merge); and (4) Stage-4: transformation and fusion. Each stage is detailed below.
Refer to caption Refer to caption Object Name Description Phrase
@@
Figure 4: Compared with applying object name (“man”), utilizing the description phrase $\mathcal{T}_{r}$ (“the man wearing a white t-shirt”) can achieve more accurate detection results.
@@
Stage-1: Object Tag Extraction. Firstly, we apply the advanced LLM, Llama3 [6], to extract key objects (e.g., “billiard table” in Fig. 3a) from the given descriptions. Each object is assigned a three-tuple form tag: $\mathcal{T}_{r}$, $\mathcal{T}_{q}$, $\mathcal{T}_{e}$}. The $\mathcal{T}_{r}$ is the object description phrase (sometimes same as name); $\mathcal{T}_{q}$ denotes the quality of object (e.g., “clear”); $\mathcal{T}_{e}$ represents the object effect on image quality (i.e., “no impact”, “positive”, or “negative”). All tag items are inferred from the description, with $\mathcal{T}_{r}$ and $\mathcal{T}_{q}$ used in later stages. The $\mathcal{T}_{e}$ item enables us to filter out non-critical objects (e.g., “image”, which refers to the whole). This explicit effect classification, similar to chain-of-thought (CoT), can reduce hallucinations. Overall, by extracting and filtering, we can obtain the target object and information for subsequent processing.
@@
Stage-2: Bounding Box Detection. Then, we detect bounding boxes for the extracted objects from the image. To accomplish this, we utilize the state-of-the-art object detection model, Grounding DINO [24]. Since multiple same-category objects may appear in one image, we utilize the $\mathcal{T}{r}$ generated Stage-1 rather than the object name for detection. For instance, in Fig. 4, the object name is “man”, and $\mathcal{T}{r}$ is “the man wearing a white t-shirt”. Leveraging “man” detects two objects (left case), while using $\mathcal{T}{r}$ can achieve the more precise result (right case).
@@
Stage-3: Box Refinement. We further refine the detected boxes. Although Stage-2 adopts $\mathcal{T}{r}$ to limit the detection range, multiple boxes may still exist. In some cases, multiple boxes may contain the wrong target. Through observations, most detection errors arise from the detection model inability to distinguish objects of same class with different quality. For instance, in Fig. 3a, for “hands”, the key (reduce image quality) is the blurry one, and the other is irrelevant. To address this problem, we design the IQA-Filter algorithm (Alg. 1). We use the MLLM-based IQA method, Q-Instruct, to verify detected bounding boxes by inputting each box patch and asking: “Is the image quality is <$\mathcal{T}_{q}$>?”, with $\mathcal{T}_{q}$ from Stage-1. We check all boxes in single-object-multiple-targets, and remove those with a “No” response.
@@
Furthermore, in some cases, multiple small or overlapping targets correspond to the same object. While these detections are accurate, an excess of targets may increase the learning difficulty for MLLMs. To address this issue, we propose the Box-Merge algorithm (Alg. 1). We merge boxes that satisfy the normalized area threshold $T_{a}$ (set to 0.256), and the overlap threshold $T_{o}$ (set to 95%). Overall, the IQA-Filter and Box-Merge algorithms effectively refine the quality of bounding boxes.
@@
$\begin{gathered}\text{idx}_{l}=y_{1}\cdot m\cdot n+x_{1}\cdot n,\\ \text{idx}_{r}=y_{2}\cdot m\cdot n+x_{2}\cdot n,\end{gathered}$ (1)
where $\text{idx}_{l}$ and $\text{idx}_{r}$ denotes the discretized coordinates. The box can be represented as $\langle\text{idx}_{l},\text{idx}_{r}\rangle$, e.g., $\langle 10,110\rangle$. Accordingly, we remap the discrete coordinates back to continuous format using the centre coordinates of each grid:
@@
where new coordinates is $\langle x_{1}^{\prime},y_{1}^{\prime},x_{2}^{\prime},y_{2}^{\prime}\rangle$. Though the discretization reduces coordinate precision, it effectively decreases the representation complexity. In our dataset, we set $n$$=$$m$$=$20, requiring at most 9 tokens (2$\times$3$+$3) for a box.
@@
1:  Input: target image $I$, object bounding boxes $\mathcal{B}$, object quality $\mathcal{T}_{q}$, area threshold $T_{a}$, overlap threshold $T_{o}$
@@
3:  Init: $\mathcal{R}\leftarrow\emptyset$ // IQA-Filter
@@
5:     $p\leftarrow\text{patch}(I,b)$; $q\leftarrow$“Is the image quality <$\mathcal{T}_{q}$>?”
6:     if Q-Instruct$(p,q)=$ ‘Yes” then
@@
[…]
@@
In conclusion, the final $\bm{answer}$ is a natural language description with precise coordinates, as shown in Fig. 3a.
For GIQA-VQA. The task requires that the $\bm{question}$ or $\bm{answer}$ relate to low-level attributes and include explicit spatial information (i.e., bounding boxes). Inspired by previous work [43, 51, 18], we apply the LLM (i.e., Llama3 [6]) to generate the corresponding QA pairs from the descriptions in GIQA-DES (depicted in Fig. 3b). We use specific templates to generate diverse QA, as follows:

### v3

# Grounding-IQA: Grounding Multimodal Language Model for Image Quality Assessment

@@ Abstract
The development of multimodal large language models (MLLMs) enables the evaluation of image quality through natural language descriptions. This advancement allows for more detailed assessments. However, these MLLM-based IQA methods primarily rely on general contextual descriptions, sometimes limiting fine-grained quality assessment. To address this limitation, we introduce a new image quality assessment (IQA) task paradigm, grounding-IQA. This paradigm integrates multimodal referring and grounding with IQA to realize more fine-grained quality perception, thereby extending existing IQA. Specifically, grounding-IQA comprises two subtasks: grounding-IQA-description (GIQA-DES) and visual question answering (GIQA-VQA). GIQA-DES involves detailed descriptions with precise locations (e.g., bounding boxes), while GIQA-VQA focuses on quality QA for local regions. To realize grounding-IQA, we construct a corresponding dataset, GIQA-160K, through our proposed automated annotation pipeline. Furthermore, we develop a well-designed benchmark, GIQA-Bench. The benchmark evaluates the grounding-IQA performance from three perspectives: description quality, VQA accuracy, and grounding precision. Experiments demonstrate that our proposed method facilitates the more fine-grained IQA application. Code: https://github.com/zhengchen1999/Grounding-IQA.
1 Introduction
@@
Image quality assessment (IQA) seeks to evaluate image quality in alignment with human perception. As a fundamental task in low-level vision, IQA is critical across multiple fields, e.g., image processing (Zhang et al., 2018; Lin et al., 2019), media transmission (Ying et al., 2020), and generative artificial intelligence (Li et al., 2023). However, this task is challenging since the human visual system is inherently subjective and complex to model (Wang et al., 2004). To enhance evaluation precision, substantial research efforts continue to be dedicated to this area (Mittal et al., 2012a; Ding et al., 2020; Wang et al., 2023; Wu et al., 2024b).
Traditional IQA methods employ handcrafted metrics to estimate quality scores (Wang et al., 2004; Mittal et al., 2012b). With advancements in deep neural networks, learning specific priors from large datasets enables more accurate score predictions (Kang et al., 2014; Bosse et al., 2017; Jinjin et al., 2020; Ke et al., 2021). Nevertheless, score-based IQA methods face challenges in complex scenarios. In such cases, image quality is influenced by multiple factors that a single score cannot effectively express (You et al., 2024c). Recently, the emergence of multimodal large language models (MLLMs) (Liu et al., 2023; Peng et al., 2024; Ye et al., 2024) with strong visual and linguistic perception capabilities provides an alternative to score-based IQA (Wu et al., 2024a; e). These models achieve more detailed and accurate image assessments through description and reasoning. However, current MLLM-based IQA methods (Wu et al., 2024b; You et al., 2024b) primarily rely on general contextual descriptions, which sometimes limits fine-grained quality assessments. For instance, in Fig. 2a, the existing method (i.e., Q-Instruct (Wu et al., 2024b)) describes the objects/areas affecting image quality through language, but cannot provide precise location information. Moreover, in Fig. 2b, for local perception, the language referring may not accurately pinpoint the target, leading to bias. These limitations restrict the application of MLLMs in comprehensive low-level perception and understanding, especially for fine-grained cases.
@@
To address these challenges and unleash the potential of MLLMs in fine-grained image quality understanding, we introduce grounding-IQA. This is a novel IQA task paradigm that integrates multimodal referring (position in) and grounding (position out) (Mao et al., 2016; Chen et al., 2023; Peng et al., 2024) with image quality assessment. This new paradigm can serve as an extension and enhancement to existing IQA methods. Specifically, we categorize grounding-IQA into two sub-tasks: (1) Grounding-IQA-Description (GIQA-DES). As illustrated in Fig. 2a, this task requires generating descriptive assessments of image quality while providing precise locations (i.e., bounding boxes) for important objects/regions impacting quality. (2) Grounding-IQA-Visual Question Answering (GIQA-VQA). As shown in Fig. 2b, this task involves QA about low-level attributes of images, especially regarding local objects. It includes addressing questions with specific coordinates (referring) or providing answers with precise positions (grounding).
@@
Since existing datasets can not realize grounding-IQA well (Liu et al., 2023; You et al., 2024a; Wu et al., 2024b), we construct a new dataset, GIQA-160K, based on the proposed paradigm. This dataset can enhance the grounding-IQA capabilities of current MLLMs. The dataset comprises 160K instruction-tuning data with 40K images from diverse domains. Specifically, the dataset corresponds to two sub-tasks: GIQA-DES includes 60K corresponding data, and GIQA-VQA contains 100K related data. To construct the corresponding dataset, we design an automated annotation pipeline. The automated pipeline generates the GIQA-160K through the public IQA dataset (Wu et al., 2024b; You et al., 2024b) (with the human-annotated description). (1) For GIQA-DES. The task includes detailed descriptions with coordinates. We generate the data through advanced vision (Liu et al., 2024c) and language (Dubey et al., 2024) models. Through these models, we extract and filter objects and corresponding coordinates from existing descriptions and images. Meanwhile, coordinates are expressed in natural language and attached to text. This avoids extra specialized tokens and ensures data compatibility. (2) For GIQA-VQA. Inspired by previous work (Wu et al., 2024b; You et al., 2024a; Li et al., 2024), we construct the required data from the detailed descriptions in GIQA-DES via the LLM. We use specific QA templates (i.e., “Yes/No”, abbreviated as Y; “What/How/Why”, abbreviated as W) and emphasize location-specific objects to generate appropriate data. The coordinates are also combined with the generated QA.
@@
Fine-tuning on the GIQA-160K dataset enables existing pre-trained MLLMs to achieve impressive grounding-IQA capabilities. As shown in Fig. 2, the fine-tuned model can ground key objects affecting image quality, and perform more fine-grained assessments based on reference coordinates. Moreover, to comprehensively evaluate the model performance on the grounding-IQA task, we propose a well-designed benchmark, GIQA-Bench. This benchmark includes 100 varying types and quality images, corresponding to 100 GIQA-Des and 150 GIQA-VQA test samples. Each sample is annotated over multiple rounds by at least three experts. We quantitatively assess grounding-IQA performance in three aspects: (1) assessment description quality (i.e., BLEU@4, LLM-Score); (2) VQA accuracy (i.e., Accuracy); and (3) grounding precision (i.e., mIoU, Tag-Recall). We test recent MLLMs, with results shown in Fig. 1. Observations indicate significant improvement in grounding-IQA after fine-tuning with GIQA-160K. Overall, our contributions are threefold:
@@ 2.1 Image Quality Assessment
Score-based Methods. Most current IQA methods are score-based. Early IQA approaches compute scores through handcrafted image data metrics (Wang et al., 2004; Moorthy & Bovik, 2011; Mittal et al., 2012a). However, these methods show a gap in quality perception compared to human judgment and are unsuitable for complex scenarios. With the development of the neural network, learning-based IQA methods have gradually become mainstream (Yang et al., 2022; Chen et al., 2024a; Shin et al., 2024). These methods leverage data-driven training to achieve more accurate quality assessments. For example, LPIPS (Zhang et al., 2018) applies the convolutional neural network to compute scores. Moreover, meta-learning (Zhu et al., 2020), multimodal models (Wang et al., 2023; Zhang et al., 2023c), and graph neural networks (Sun et al., 2022) have been adopted to further improve IQA. However, score-based IQA methods face limitations in complex scenarios. The simple score cannot effectively represent the multiple aspects affecting image quality.
MLLM-based Methods. Multimodal large language models (MLLMs) exhibit remarkable multimodal (language/vision) understanding by integrating visual modules into LLMs (Liu et al., 2023; Zhang et al., 2023a; Ye et al., 2024). MLLMs achieve outstanding performance in various multimodal tasks, including visual question answering and image captioning. Recently, several studies have also demonstrated the potential of MLLMs in low-level visual perception and assessment (Wu et al., 2024b; You et al., 2024b; b; Wu et al., 2024d; Chen et al., 2024b). For instance, Q-Instruct (Wu et al., 2024b) constructs a multimodal dataset to enhance. Q-Align (Wu et al., 2024c) guides MLLMs in scoring by defining discrete text-based levels. DepictQA (You et al., 2024c) enables quality comparison and reasoning based on reference images. These approaches advance the application of MLLMs in IQA, achieving more accurate assessments. Nevertheless, these models primarily rely on contextual descriptions, and face limitations in fine-grained applications, e.g., local perception.
@@ 2.2 Multimodal Referring and Grounding
Multimodal spatial perception involves referring and grounding. Referring requires the model to understand the specific region based on position input, e.g., region-level captioning (Krahmer & Van Deemter, 2012; Zellers et al., 2019). Grounding, on the other hand, involves the model describing the region by outputting position, e.g., referring expression comprehension (Kazemzadeh et al., 2014; Luo & Shakhnarovich, 2017). Currently, MLLMs perform impressively in spatial perception, further advancing these tasks. Some methods focus on grounding, achieving complex reasoning (Lai et al., 2024) or multi-object (Ren et al., 2024) segmentation. Meanwhile, other approaches, e.g., GPT4RoI (Zhang et al., 2023b), emphasize understanding specific regions (referring). Furthermore, some works unify referring and grounding (Chen et al., 2023; Li et al., 2024; Rasheed et al., 2024; Peng et al., 2024; You et al., 2024a). Additionally, in IQA, Q-Ground (Chen et al., 2024b) achieves degradation region grounding but lacks referring capabilities. In contrast, our Grounding-IQA integrates multimodal referring and grounding with IQA to enhance quality perception.
Figure 3: The illustration of the automated annotation pipeline. (a) GIQA-DES Pipeline: Constructs the $\bm{answer}$ from the given image and description via a four-stage process, while the $\bm{question}$ comes from a predefined question pool. (b) GIQA-VQA Pipeline: Generates the corresponding QA data utilizing descriptions from GIQA-DES and the LLM (Llama3 (Dubey et al., 2024)).
@@ 3.1 Grounding-IQA
GIQA-DES. The task requires the model to provide a detailed description of image quality. Additionally, it needs accurate location information (e.g., bounding box) for key objects/regions that impact image quality, as shown in Fig. 5a. This corresponds to the fact that humans consider not only the overall quality (e.g., image clarity) but also the quality of specific objects or locations when assessing image quality. Meanwhile, accurate location information also enables targeted information for downstream tasks (e.g., image editing). This task is similar to grounded image captioning (Zhou et al., 2020), but places greater emphasis on low-level attributes. While some MLLMs (Chen et al., 2023; Peng et al., 2024; Li et al., 2024) perform well in grounded image captioning, they still struggle with quality perception. We demonstrated it in Sec. 4.3.
GIQA-VQA. The second task focuses on the question-answering ability in low-level perception, particularly for local objects. Corresponding to multimodal referring and grounding, this task can be divided into two scenarios. Referring: querying low-level attributes in the specified region (input position), as shown in Fig. 5b. Grounding: providing answers that include specific locations (output position) based on the question, as depicted in Fig. 5b. These two scenarios are related to region captioning (Zhou et al., 2020) and phrase grounding (Zhou et al., 2020), respectively. However, like GIQA-DES, GIQA-VQA involves quality perception, which is challenging for current MLLMs.
@@ 3.2 Automated Annotation Pipeline
Data is essential for achieving Grounding-IQA. Therefore, we construct an automated annotation pipeline to generate data (i.e., GIQA-160K). This pipeline leverages public IQA datasets (Wu et al., 2024b; You et al., 2024b) that contain human-annotated descriptions. Following previous schemes (Liu et al., 2023; Ye et al., 2024), the data format is $\bm{image}$, $\bm{question}$, $\bm{answer}$}. The $\bm{image}$ is the evaluation target. Depending on the sub-task, the $\bm{question}$ and $\bm{answer}$ fields may include precise coordinates (i.e., bounding box), in addition to text. The illustration of the whole pipeline is in Fig. 3. Besides, more details are provided in the supplementary material.
Object Name Description Phrase
@@
Figure 4: Utilizing the description phrase $\mathcal{T}_{r}$ (“the man wearing a white t-shirt”) yields more accurate detection than applying object name (“man”).
For GIQA-DES. In this task, the $\bm{question}$ is relatively fixed, as in Fig. 5a. For each data point, the $\bm{question}$ is randomly selected from the question pool with 15 similar questions. For the $\bm{answer}$, it is a detailed description with coordinates. We construct it via a four-stage process from existing images and associated description, as illustrated in Fig. 3: (1) Stage-1: object tag extraction; (2) Stage-2: bounding box detection; (3) Stage-3: box refinement (filter and merge); and (4) Stage-4: transformation and fusion. Each stage is detailed below.
@@
Stage-1: Object Tag Extraction. Firstly, we apply the advanced LLM, i.e., Llama3 (Dubey et al., 2024), to extract key objects (e.g., “billiard table” in Fig. 3a) from the given descriptions. Each object is assigned a three-tuple form tag: $\mathcal{T}_{r}$, $\mathcal{T}_{q}$, $\mathcal{T}_{e}$}. The $\mathcal{T}_{r}$ is the object description phrase (sometimes same as name); $\mathcal{T}_{q}$ denotes the quality of object (e.g., “clear”); $\mathcal{T}_{e}$ represents the object effect on image quality (i.e., “no impact”, “positive”, or “negative”). All tag items are inferred from the description, with $\mathcal{T}_{r}$ and $\mathcal{T}_{q}$ used in later stages. The $\mathcal{T}_{e}$ item enables us to filter out non-critical objects (e.g., “image”, which refers to the whole). This explicit effect classification, similar to chain-of-thought (CoT), can reduce hallucinations.
@@
Stage-2: Bounding Box Detection. Then, we detect bounding boxes for the extracted objects from the image. To accomplish this, we utilize the state-of-the-art object detection model, Grounding DINO (Liu et al., 2024c). Since multiple same-category objects may appear in one image, we utilize the $\mathcal{T}{r}$ generated Stage-1 rather than the object name for detection. For instance, in Fig. 4, the object name is “man”, and $\mathcal{T}{r}$ is “the man wearing a white t-shirt”. Leveraging “man” detects two objects (left case), while using $\mathcal{T}{r}$ can achieve the more precise result (right case).
@@
Stage-3: Box Refinement. Although Stage-2 adopts $\mathcal{T}{r}$ to limit the detection range, multiple boxes may still exist. In some cases, multiple boxes may contain the wrong target. Through observations, most detection errors arise from the detection model inability to distinguish objects of same class with different quality. For instance, in Fig. 3a, for “hands”, the key (reduce image quality) is the blurry one, and the other is irrelevant. To address this problem, we design the IQA-Filter algorithm (Alg. 1). We use the MLLM-based IQA method, Q-Instruct, to verify detected bounding boxes by inputting each box patch and asking: “Is the image quality is $\mathcal{T}_{q}$>?”, with $\mathcal{T}_{q}$ from Stage-1. We check all boxes in single-object-multiple-targets, and remove those with a “No” response.
@@
Furthermore, in some cases, multiple small or overlapping targets correspond to the same object. While these detections are accurate, an excess of targets may increase the learning difficulty for MLLMs. To address this issue, we propose the Box-Merge algorithm (Alg. 1). We merge boxes that satisfy the normalized area threshold $T_{a}$ (set to 0.256), and the overlap threshold $T_{o}$ (set to 95%).
@@
$\begin{gathered}\text{idx}_{l}=y_{1}\cdot m\cdot n+x_{1}\cdot n,\quad\text{idx}_{r}=y_{2}\cdot m\cdot n+x_{2}\cdot n,\end{gathered}$ (1)
where $\text{idx}_{l}$ and $\text{idx}_{r}$ denotes the coordinates. The box can be represented as $\langle\text{idx}_{l},\text{idx}_{r}\rangle$. Accordingly, we remap the discrete coordinates back to a continuous format using the centre coordinates:
@@
where new coordinates is $\langle x_{1}^{\prime},y_{1}^{\prime},x_{2}^{\prime},y_{2}^{\prime}\rangle$. Though the discretization reduces coordinate precision, it effectively simplifies the representation. In our dataset, we set nn==mm==20, requiring at most 9 tokens.
@@
1:Input: target image II, object bounding boxes $\mathcal{B}$, object quality $\mathcal{T}_{q}$, area threshold $T_{a}$, overlap threshold $T_{o}$
@@
3:Init: $\mathcal{R}\leftarrow\emptyset$ $\triangleright$ IQA-Filter: filter boxes by quality query
@@
5:  $p\leftarrow\text{patch}(I,b)$; $q\leftarrow$“Is the image quality $\mathcal{T}_{q}$>?”
6:  if Q-Instruct(p,q)=(p,q)= ‘Yes” then
@@
13:   if $\text{area}(\mathcal{R}[i])<T_{a}$ and $\text{is-touch}(\mathcal{R}[i],\mathcal{R}[j])$ or $\text{coverage-ratio}(\mathcal{R}[i],\mathcal{R}[j])>T_{o}$ then
@@
Finally, the $\bm{answer}$ is a natural language description with precise coordinates, as shown in Fig. 3a.
For GIQA-VQA. The task requires that the $\bm{question}$ or $\bm{answer}$ relate to low-level attributes and include explicit spatial information (i.e., bounding boxes). Inspired by previous work (Wu et al., 2024b; You et al., 2024a; Li et al., 2024), we apply the LLM (i.e., Llama3 (Dubey et al., 2024)) to generate the corresponding QA pairs from the descriptions in GIQA-DES (depicted in Fig. 3b). We use specific templates to generate diverse QA. Details are as follows: